OpenStudy (twizttiez):

What is the point-slope form of the equation of a line that passes through the point (3, 5) and has a slope of 2? A.y+3=2(x+5) B.y-3=2(x-5) C.y-5=2(x-3) D.y-5=-2(x-3)

OpenStudy (twizttiez):

@dmndlife24 @silverr @dimensionx

OpenStudy (dmndlife24):

C.

OpenStudy (twizttiez):

Ok 100% sure?

OpenStudy (anonymous):

its c

OpenStudy (anonymous):

i checked on my test

OpenStudy (dmndlife24):

y - (5) = m (x - (3)) y - 5 = 2 (x - 3) Yes 100% sure :D

OpenStudy (twizttiez):

Ok thank you both i have two more would you mind helping?

OpenStudy (anonymous):

c seems accurate.

OpenStudy (twizttiez):

I mean 3

OpenStudy (twizttiez):

Do you want me to open a new window?